OpenStudy (anonymous):

Please Help! (New to Physics) WILL Medal and fan! Do a dimensional analysis of the equation below. (I will add problem in proper format) 1/2 m v^2 = 1/2 m v base 0^2 +sqrt(mgh)

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }mv ^{2} = \frac{ 1 }{ 2 }mv _{0}^{2} + \sqrt{mgh}\]

OpenStudy (irishboy123):